(1) Let Q(x, y) be the statement "x+y=x-y." If the domain for both variables consists of all integers, what are the truth values?
(a)Q(1,1)
(b)Q(2,0)
(c)∀yQ(1, y)
(d)∃xQ(x,2)
(e)∃x∃yQ(x, y)
(f)∀x∃yQ(x, y)
(g)∃y∀xQ(x, y)
(h)∀y∃xQ(x, y)
(i)∀x∀yQ(x, y)
1
Expert's answer
2020-09-23T17:38:24-0400

Solution to a:

Q(1,1)


"x+y=1+1=2""x-y=1-1=0"

since x+y is not equal to x-y, the truth value of Q(x,y) is FALSE

Ans: FALSE


Solution to b:

Q(2,0)


"x+y=2+0=2""x-y=2-0=2"

since x+y=x-y, the truth value of Q(x,y) is TRUE.

Ans: TRUE


Solution to c:

"\forall y Q(1,y)"

let y=1;


"x+y=1+1=2""x-y=1-1=0"

since "\exists y" for which x+y is not equal to x-y (e.g y=1), the truth value of Q(x,y) is FALSE.

Ans: FALSE


Solution to d:

"\exists xQ(x,2)"


"x+2=x-2""x-x=-2-2""0=-4"

since the above equation does not give a definite solution of x, the truth value of Q(x,y) is FALSE.

Ans: FALSE


Solution to e:

"\exists x\exists yQ(x,y)"

let x=2, y=0;


"x+y=2+0=2""x-y=2-0=2"

since "\exists x" and "\exists y" such that x+y=x-y (i.e x=2, y=0), the truth value of Q(x,y) is TRUE.

Ans: TRUE


Solution to f:

"\forall x \exists y Q(x,y)"

let x=1,y=1;


"x+y=1+1=2""x-y=1-1=0"

since x+y is not equal x-y when x=1 and y=1, the truth value of Q(x,y) is FALSE

Ans: FALSE


Solution to g:

"\exists y \forall xQ(x,y)"

when y=0;

"x+0=x-0""x=x"

since x+y=x-y for all x when y=0, the truth value of Q(x,y) is TRUE

Ans: TRUE


Solution to h:

"\forall y \exists x Q(x,y)"

let y=1;


"x+1=x-1""x-x=-1-1""0=-2"

since the above equation does not give a definite solution of x, the truth value of Q(x,y) is FALSE.

Ans: FALSE


Solution to i:

"\forall x\forall yQ(x,y)"

let x=1, y=1;


"x+y=1+1=2""x-y=1-1=0"

since x+y is not equal to x-y, the truth value of Q(x,y) is FALSE

Ans: FALSE